# Spineless Traversal for Layout Invalidation

### v4

@@ Abstract.
Latency is a major concern for web rendering engines like those in Chrome, Safari, and Firefox. These engines reduce latency by using an incremental layout algorithm to redraw the page when the user interacts with it. In such an algorithm, elements that change frame-to-frame are marked dirty, and only the dirty elements need be processed to draw the next frame, dramatically reducing latency. However, the standard incremental layout algorithm must search the page for dirty elements, accessing a number of auxiliary elements in the process. These auxiliary elements add cache misses and stalled cycles, and are responsible for a sizable fraction of all layout latency.
@@
Web browsers, incremental, order maintenance, latency
@@ 1. Introduction
Latency is a major concern for modern web rendering engines such as those in Chrome, Safari, and Firefox. A rendering engine must redraw pages 60 times per second to guarantee smooth animations, fluid interactions, and responsive applications as users browse and interact with web pages. When this frame rate cannot be met, the user experiences lag and may be forced to use another web application, browser, or device. Modern 120 Hz displays demand even lower latency.
Layout is a key driver of web rendering latency. Layout means calculating the size and position of each element of the web page, after which the page can be rendered into pixels on the screen. Every time the user interacts with the web page by hovering over an element, receiving updated data, or even observing an animation, the web page changes and must be re-laid-out. Since layout is only one part of the larger rendering pipeline this must be completed in a millisecond or less in order to meet the 60 frame-per-second goal. On such a tight budget, every cycle counts!
@@
The key optimization that makes this possible is incrementalization. When an element on the page changes, the browser marks it as dirty. When the page is re-laid-out, the rendering engine traverses the page tree to find and re-lay-out only the dirty nodes. In the process, it might mark additional nodes dirty, due to dependencies between nodes, but typically—in, say, animations or small interactions—only a few nodes are ultimately re-laid-out. In these cases, searching the tree for dirty elements is the bottleneck, especially since every node access is likely to incur a cache miss.
Browsers thus use the “Double Dirty Bit” algorithm to find dirty nodes more quickly (Panchekha and Harrelson, 2024). This algorithm adds summary bits to each node to identify and skip subtrees without any dirty nodes. While this reduces the search time, Double Dirty Bit still has to traverse the tree, starting from the root, to find dirty nodes, and must access not only the actually-dirty nodes but many extra “auxiliary nodes”. On large pages, auxiliary nodes can significantly outnumber the actually-dirty nodes; since each node access can cause cache miss, simply traversing these auxiliary nodes, just to check their summary bits, stalls the layout algorithm for hundreds of microseconds. Indeed, this problem is widely observed in practice; Google’s widely-used web performance tool, Lighthouse, uses tree depth and maximum children count as metrics precisely because these parameters determine the number of auxiliary nodes.
@@
We introduce Spineless Traversal: a new, faster algorithm for incremental layout. Unlike the standard Double Dirty Bit algorithm, Spineless Traversal accesses only dirty nodes; it requires no auxiliary nodes. It therefore reduces cache misses. Spineless Traversal works by storing the set of dirtied nodes in a queue and jumping directly from one dirty node to the next, with no auxiliary nodes in between.
The key challenge is traversing the dirty nodes in the correct order. Recomputing a field on one node can mark fields on other nodes dirty, and the set of transitive dependencies is complex. Fields must therefore be recomputed in a specific order, and Spineless Traversal must respect that order as it jumps from node to node. Spineless Traversal thus applies a timestamp to each node, and uses a priority queue traverse nodes in timestamp order. To maintain the order as nodes are added and removed, Spineless Traversal uses an order maintenance data structure to compute relative timestamps in a flexible way. Both the priority queue and order maintenance structure are heavily optimized to make them competitive with Double Dirty Bit.
@@
We implement Spineless Traversal in Megatron, a new compiler for incremental layout algorithms. Megatron implements decades of research on attribute grammars: it automatically analyzes the static dependency graph, synthesizes recomputation and marking functions, and guarantees a correct incremental layout. It implements standard optimizations (unboxing, interning, field packing, and jump tables) and compiles layout algorithms to highly efficient C++ code. Megatron supports both Double Dirty Bit and Spineless Traversal via a common invalidation traversal interface, allowing an apples-to-apples comparison between them.
We compare the two algorithms on a significant fragment of web layout that includes line breaking, flex-box, intrinsic sizes, and many other features, using 50 of real-world web pages like Twitter, Discord, Github and Lichess. Across 2216 frames, Spineless Traversal is 1.78×1.78$\times$1.78 × times faster on average. Speedups are concentrated in the most latency-critical changes with the fewest dirty nodes: on the 60.7% of benchmarks where fewer than 1% of fields are recomputed, Spineless Traversal achieves a speedup of 2.35×2.35$\times$2.35 × or more.
@@ 2. Web Layout Background
Complex web (Facebook, Amazon, GMail), desktop (VS Code, Figma, Slack), and mobile (WeChat) applications are driven by the interaction of two software components: the application itself and the browser. The application code, typically in JavaScript, implements the actual application logic and interacts with the browser by making DOM API calls that modify the browser’s internal representation of the HTML and CSS that defines the web page. The browser code then executes its “rendering pipeline”—event handling, hit testing, matching, styling, layout, paint, layerizing, rastering, and drawing—which reads this internal representation and transforms it, step by step, into pixels on the screen. Application code is typically most of the execution time but that code’s effects are ultimately visible to the user only through “system calls” serviced by browser-level code. For the application to be responsive and smooth, those calls have to be serviced in 16 milliseconds (60 frames per second).
Concretely, a browser application is written in HTML, which defines a tree called the “document”, whose nodes include text, buttons, and structural elements like divs. The application code binds specific callback functions to user actions like clicking on certain elements, [^1] Callbacks can also run in response to timers, network requests, or a dizzying array of other events. and these callbacks can modify the HTML using DOM methods like appendChild or setAttribute. When the callback finishes executing, [^2] Or after multiple callbacks finish executing, as decided by the browser’s task scheduler. the browser executes the full rendering pipeline to display the results to the user. Some DOM methods like getBoundingRect or offsetX read state computed during the rendering pipeline; calling these methods can requires the browser to perform additional passes through the rendering pipeline, and is necessary for common interactions like tooltips.
@@ 2.1. The Layout Phase
The layout tree is basically the HTML tree, with minor differences like “generated content” (like bullets for list items) and “fragmentation” (like line breaking) that are not critical to this paper. The trees are both big and unbalanced; the famously minimal Google home page page, for example, has 842 nodes, with a maximum fanout of 16 and depth of 18; it is drawn in Figure 1. In memory, the layout tree is stored as a pointer tree, with the children of each node stored in a doubly-linked list.333 This ensures that node insertions and deletions are fast, even in poorly-balanced trees. The application can add or remove nodes from this tree, or write to their “properties” and “attributes”, to update what the user sees.
@@
Each node’s layout fields depend on the layout fields of its neighbors. For example, imagine a paragraph containing several lines; both the paragraph and each line would correspond to a layout node. The height of the paragraph, in this case, would be the sum of the heights of all its children plus any gaps between them, while a node’s width would be its parent’s width, minus some padding. Of course, real-world web layout is much more complex; Chrome’s implementation is about 100,000 lines of code and computes hundreds of fields per node. To compute all these fields correctly, the layout algorithm recursively traverses the tree multiple times. For example, the layout algorithm might first compute the intrinsic size of each element in a bottom-up traversal; then compute preferred sizes, in a top-down traversal; and then apply flexible sizing rules in a bottom-up traversal to finally compute each element’s actual size. Note that not only are multiple passes necessary, but that each pass must visit the nodes of the layout tree in a specific order so that each layout field’s dependencies are satisfied.
@@
Figure 2. A trace of nine milliseconds of Chrome opening Youtube; time flows left to right while the call stack grows down. The violet “recalculate style” and “layout” blocks are phases of the browser rendering pipeline, with layout in total consuming four milliseconds; the faint vertical lines are half-millisecond marks. Layout happens multiple times in one frame due to “forced reflows”. The “Summary” tab at the bottom shows that the selected (longest) layout only updates 23 of 2 01 nodes in the layout tree.
@@
Figure 2 shows layout performance in practice: a trace of Google Chrome opening YouTube, captured and visualized using Chrome’s “Performance” tools. The trace covers 9 milliseconds of execution time, with time running left to right. Function calls grow downward, and the leaves, labeled “recalculate style”, “layout”, and “pre-paint” (barely visible) are different phases of the rendering pipeline; the application-level code is offscreen to the left. In these nine milliseconds, the application requests layout four times due to calling APIs that require (“forced reflow”) multiple passes through the rendering pipeline. These four layout each take more time than the browser developers’ target latency, as indicated by the red tags on these blocks.
@@
Below the trace, the “Summary” tab shows more information about a single, selected layout taking 2.1 milliseconds. This layout operated on a layout tree of 2 301 nodes, starting from the root of the web page (#document). However, it only needed to update 23 of them; in other words, this layout was incremental. The long running time of the layout phase nonetheless caused unacceptable latency; in fact, this frame “drops”, meaning the browser isn’t able to update the page in time to show smooth animations and interactions.
@@ 2.2. Formal Modeling
Luckily, the programming languages community has developed a sophisticated understanding of layout. Early work by Meyerovich and others (Jones et al., [n. d.]; Meyerovich and Bodik, 2010; Meyerovich, 2013) developed “attribute grammars” as a formalism for defining layout rules and implementations. The later Cassius project (Panchekha and Torlak, 2016; Panchekha et al., 2018; Prazina et al., 2023) developed a full, standards-compliant implementation of a significant fragment of the layout rules in this formalism, and the Hecate and Medea tools (Chen et al., 2022; Liu et al., 2023) have shown that automatic synthesis can be scaled to such fragments.
@@
Figure 3 defines an attribute grammar. In an attribute grammar, a layout is defined by a set of passes (the rules) performed in a certain order (the schedule). Each pass performs a recursive, in-order traversal of the tree, computing some fields pre-order and some fields post-order; every field is written to exactly once in exactly one pass. For each field assignment $\mathsf{self}.V\leftarrow T$, the expression $T$ can refer to fields of $\mathsf{self}$ or to fields of its $\mathsf{parent}$, $\mathsf{prev}$ and $\mathsf{next}$ sibling, or $\mathsf{first}$ and $\mathsf{last}$ child; expressions can also test whether a given neighbor exists ($\mathsf{N?}$). Computations can also refer to attributes or properties of the current node using $\mathsf{attribute}[x]$ or $\mathsf{property}[x]$.444 The split HTML attributes and CSS properties use two different namespaces because some names, like height, appear in both sets. There is no other semantic difference between them, and other accessible properties, such as the tag name or image width and height are modeled in our system as special properties. All computations besides field assignments are pure, there are no other loops or data structures, and the only field access allowed is to a node’s neighbors. Despite this, even complex layout features like flexible box layout are expressible in such a DSL.
@@
Figure 5 shows the layout example for paragraphs and lines, defined using an attribute grammar. Here the nodes are the paragraphs and lines, and they all have three layout fields: a width $W$, a height $H$, and an intermediate field $HA$ that computes the height of a node and its previous siblings (plus any gaps between lines). Width information propagates from parents to children, starting at 50 pixels and subtracting, at each level, 5 pixels of padding on the left and right. Height information propagates from children to parents—a node’s height is the sum of its children’s heights, plus 5-pixel gaps between lines—and relies on the intermediate $HA$ field which propagates information from previous to next siblings. Note that the idea of “summing all children’s heights” is not expressed with a loop; instead, it is expressed as an additional layout field, which has its own computation rule.
@@
A key property of attribute grammar DSLs like this one is that dependencies and traversal orders are static. Specifically, for any field assignment, examining the expression $T$ reveals which other fields on which other nodes it depends on. These static dependencies are critical for invalidation. Also, the schedule language ensures that the relative order in which two fields are computed never changes, even as nodes are added or removed. This is critical for Spineless Traversal. In this paper, we assume that this schedule respects the field dependencies and that the rules have no cyclic dependencies; we also assume that the schedule has already been optimized by fusing traversals and ordering field assignments. A substantial literature exists on these topics (grafter; Chen et al., 2022; Liu et al., 2023), and the layout implementation used in our evaluation is detailed in Section 7.1. In any case, while this paper focuses on web layout, we expect Spineless Traversal to be applicable to incremental computations in other domains, including in compilation, static analysis, computer graphics, and databases, as long as they can be expressed as an attribute grammar in this or a similar DSL.
@@ 3. Incremental Layout
An incremental layout algorithm reuses layout field values between layouts as the layout tree is modified by the application code. For example, when the user moves their mouse, a tooltip may need to move to follow the cursor; incremental layout would change the position of the tooltip while reusing all other layout fields for all other elements. Most layouts change only a small fraction of fields, especially the most latency-critical interactions like hovers, drags, animations, and text editing, so incremental layout can dramatically faster than computing each layout from scratch.

### v5

@@ Abstract.
Latency is a major concern for web rendering engines like those in Chrome, Safari, and Firefox. These engines reduce latency by using an incremental layout algorithm to redraw the page when the user interacts with it. In such an algorithm, elements that change frame-to-frame are marked dirty, and only those elements are processed to draw the next frame, dramatically reducing latency. However, the standard incremental layout algorithm must search the page for dirty elements, accessing auxiliary elements in the process. These auxiliary elements add cache misses and stalled cycles, and are responsible for a sizable fraction of all layout latency.
@@
Web Browsers, Layout, Incremental Computing, Order Maintenance, Latency
copyright: ccdoi: 10.1145/3729322journalyear: 2025journal: PACMPLjournalvolume: 9journalnumber: PLDIarticle: 219publicationmonth: 6ccs: Software and its engineering Source code generationccs: Software and its engineering Translator writing systems and compiler generatorsccs: Software and its engineering Domain specific languagesccs: Theory of computation Design and analysis of algorithms
@@ 1. Introduction
Latency is a major concern for modern web rendering engines. A rendering engine, such as that in Chrome, Firefox, or Safari, must redraw pages 60 times per second to guarantee smooth animations, fluid interactions, and responsive web applications. When this frame rate cannot be met, the user experiences lag and may be forced to use another web application, browser, or device. Modern 120 Hz displays demand even lower latency.
Layout is a key driver of web rendering latency. Layout means calculating the size and position of each element on the web page, after which the page can be rendered as pixels on the screen. Every time the user interacts with the web page by hovering over an element, receiving updated data, or even observing an animation, the web page changes and must be re-laid-out. Since layout is only one part of the larger rendering pipeline, this re-layout must be completed in a millisecond or less in order to meet the 60 frame-per-second goal. On such a tight budget, every cycle counts!
@@
The key optimization that makes this possible is incrementalization. When an element on the page changes, the browser marks it dirty. When the page is re-laid-out, the rendering engine traverses the page to find and re-lay-out only the dirty elements. That might mark additional elements dirty, due to dependencies between elements, but typically—in, say, animations—only a few nodes are ultimately re-laid-out. In these cases, searching the tree for dirty elements is the bottleneck, especially since every element access is likely to incur a cache miss.
Browsers use the “Double Dirty Bit” algorithm to find dirty nodes more quickly (Panchekha and Harrelson, 2024; Garseil, [n. d.]). This algorithm adds summary bits to identify and skip subtrees without any dirty elements. While this reduces the search time, Double Dirty Bit still has to traverse the tree, starting from the root, to find dirty elements, and thus accesses not only the actually-dirty elements but many extra “auxiliary nodes”. On large pages, auxiliary nodes can significantly outnumber the actually-dirty elements; and since each node access can cause a cache miss, simply traversing these auxiliary nodes, just to check their summary bits, can stall the layout algorithm for hundreds of microseconds. This problem is widely observed in practice; Google’s widely-used web performance tool, Lighthouse, measures tree depth and maximum children count precisely because these parameters determine the number of auxiliary nodes.
@@
We introduce Spineless Traversal: a new, faster algorithm for incremental layout. Unlike Double Dirty Bit, Spineless Traversal accesses only dirty elements, not auxiliary nodes. It therefore reduces cache misses. Spineless Traversal works by storing the set of dirty elements in a queue and jumping directly from one to the next, with no auxiliary nodes in between.
The key challenge is traversing the dirty nodes in the correct order. Recomputing a field on one node can mark fields on other nodes dirty, and the set of transitive dependencies is complex. Fields must therefore be recomputed in a specific order, and Spineless Traversal must respect that order as it jumps from node to node. Spineless Traversal thus stores a timestamp on each node, and uses a priority queue to traverse nodes in timestamp order. To maintain timestamps as nodes are added and removed, Spineless Traversal order maintenance, which compute relative timestamps in a flexible way. Both the priority queue and order maintenance structure are heavily optimized to make them competitive with Double Dirty Bit.
@@
We implement Spineless Traversal in Megatron, a new compiler for incremental layout algorithms. Megatron implements decades of research on attribute grammars: it statically analyzes the dependency graph, synthesizes recomputation and marking functions, and guarantees a correct incremental layout. It implements standard optimizations (unboxing, interning, field packing, and jump tables) and compiles layout algorithms to highly efficient C++ code. Megatron supports both Double Dirty Bit and Spineless Traversal via a common invalidation traversal interface, allowing an apples-to-apples comparison between them.
We compare the two algorithms on a significant fragment of web layout that includes line breaking, flex-box, intrinsic sizes, and many other features, benchmarking 50 real-world web pages like Twitter, Discord, Github, and Lichess. Across 2216 frames, Spineless Traversal is 1.80×1.80$\times$1.80 × times faster on average. Speedups are concentrated in the most latency-critical frames: on the 65.6% of frames where at most 1% of fields are recomputed, Spineless Traversal achieves a speedup of $2\times2$.
@@ 2. Web Layout Background
Complex web (Facebook, Amazon, GMail), desktop (VS Code, Figma, Slack), and mobile (WeChat) applications involve two software components: the application itself and the browser. The application-level code, typically in JavaScript, implements the actual application logic and interacts with the browser by making DOM API calls that modify the browser’s internal representation of the web page. The browser then executes its “rendering pipeline”—event handling, hit testing, matching, styling, layout, paint, layerizing, rastering, and drawing—which reads this internal representation and transforms it, step by step, into pixels on the screen. Most execution time is typically spent in application code, but that code’s effects are ultimately visible to the user only through “system calls” serviced by browser-level code. For the application to be responsive and smooth, those calls have to be serviced in 16 milliseconds (60 frames per second).
Concretely, a browser application is written in HTML, which defines a tree called the “document”, whose nodes include text, buttons, and structural elements like divs. The application code binds specific callback functions to user actions like clicking on a button, [^1] Callbacks can also run in response to timers, network requests, or a dizzying array of other events. and these callbacks can modify the HTML using DOM methods like appendChild or setAttribute. When the callback finishes executing, [^2] Or after multiple callbacks finish executing, as decided by the browser’s task scheduler. the browser executes the full rendering pipeline to display the results to the user. Also, some DOM methods like getBoundingRect or offsetX read state computed during the rendering pipeline; calling these methods can requires the browser to perform additional passes through the rendering pipeline. This is necessary for some common interactions like tooltips.
@@ 2.1. The Layout Phase
The layout tree is basically the HTML tree, with minor differences for “generated content” (like bullets for list items) and “fragmentation” (like line breaking) that are not critical to this paper. The trees are both big and unbalanced; the famously minimal Google home page page, for example, has 842 nodes, with a maximum fanout of 16 and depth of 18; it is drawn in Figure 1. In memory, the layout tree is stored as a pointer tree, with the children of each node stored in a doubly-linked list.333 This ensures that node insertions and deletions are fast, even in poorly-balanced trees. The application can add or remove nodes from this tree, or write to their “properties” and “attributes”, to update what the user sees.
@@
Each node’s layout fields depend on the layout fields of its neighbors. For example, imagine a paragraph containing several lines; there would be a layout node for the paragraph and another for each line as children of the paragraph’s. The height of the paragraph, in this case, would be the sum of the heights of all its children plus any gaps between them, while a line’s width would be its parent’s width, minus some padding. Of course, real-world web layout is much more complex; Chrome’s implementation is about 100,000 lines of code and computes hundreds of fields per node. To compute all these fields correctly, the layout algorithm recursively traverses the tree multiple times. For example, the layout algorithm might first compute the intrinsic size of each element in a bottom-up traversal; then compute preferred sizes, in a top-down traversal; and then apply flexible sizing rules in a bottom-up traversal to finally compute each element’s actual size. Note that not only are multiple passes necessary, but that each pass must visit the nodes of the layout tree in a specific order so that each layout field’s dependencies are satisfied.
@@
Figure 2. A trace of nine milliseconds of Chrome opening Youtube; time flows left to right while the call stack grows down. The violet “recalculate style” and “layout” blocks are phases of the browser rendering pipeline, with layout in total consuming four milliseconds; the faint vertical lines are half-millisecond marks. Layout happens multiple times in one frame due to “forced reflows”. The “Summary” tab at the bottom shows that the selected (longest) layout only updates 23 of 2 301 nodes in the layout tree.
@@
Figure 2 shows layout performance in practice: a trace of Google Chrome opening YouTube, captured and visualized using Chrome’s “Performance” tools. The trace covers 9 milliseconds of execution time, with time running left to right. Function calls grow downward, and the leaves, labeled “recalculate style”, “layout”, and “pre-paint” (barely visible) are different phases of the rendering pipeline; the application-level code is offscreen to the left. In these nine milliseconds, the application requests layout four times by calling APIs that require (“forced reflow”) multiple passes through the rendering pipeline. These four layouts each take longer than the browser developers’ target latency.
@@
Below the trace, the “Summary” tab shows more information about a single, selected layout taking 2.1 milliseconds. This layout operates on a layout tree of 2 301 nodes, starting from the root of the web page (#document). However, it only needed to update 23 of them; in other words, this layout was incremental. The long running time of the layout phase nonetheless caused unacceptable latency; in fact, this frame “drops”, meaning the browser isn’t able to update the page in time to show smooth animations and interactions.
@@ 2.2. Formal Modeling
Luckily, the programming languages community has developed a sophisticated understanding of layout. Early work by Meyerovich and others (Jones et al., [n. d.]; Meyerovich and Bodik, 2010; Meyerovich, 2013) developed “attribute grammars” as a formalism for defining layout rules and implementations. The later Cassius project (Panchekha and Torlak, 2016; Panchekha et al., 2018; Prazina et al., 2023) developed a full, standards-compliant implementation of a significant fragment of web layout in this formalism, and the Hecate and Medea tools (Chen et al., 2022; Liu et al., 2023) have shown that automatic synthesis can be scaled to such fragments.
@@
Figure 3 defines an attribute grammar. An attribute grammar is defined by a set of passes (the rules) performed in a certain order (the schedule). Each pass performs a recursive, in-order traversal of the tree, computing some fields pre-order and some fields post-order; every field is written to exactly once in exactly one pass. For each field assignment $\mathsf{self}.V\leftarrow T$, the expression $T$ can refer to fields of $\mathsf{self}$ or to fields of its $\mathsf{parent}$, $\mathsf{prev}$ and $\mathsf{next}$ sibling, or $\mathsf{first}$ and $\mathsf{last}$ child; expressions can also test whether a given neighbor exists ($\mathsf{N?}$). Computations can also refer to attributes or properties of the current node using $\mathsf{attribute}[x]$ or $\mathsf{property}[x]$.444 HTML attributes and CSS properties use two different namespaces because some names, like height, appear in both sets; there is no other semantic difference. Other accessible properties, such as the tag name or image width and height are modeled in our implementation as special properties. All computations besides field assignments are pure, there are no other loops or data structures, and the only field access allowed is to a node’s neighbors. Despite this, even complex layout features like flexible box layout are expressible in such a DSL.
@@
Figure 5 shows example layout rules for paragraphs and lines, defined using an attribute grammar. Here the nodes for paragraphs and lines have three layout fields: a width $W$, a height $H$, and an intermediate field $HA$ that computes the height of a node and its previous siblings (plus any gaps between lines). Width information propagates from parents to children, starting at 50 pixels and subtracting, at each level, 5 pixels of padding on the left and right. Height information propagates from children to parents—a node’s height is the sum of its children’s heights, plus 5-pixel gaps between lines—and relies on the intermediate $HA$ field which propagates information from previous to next siblings. Note that the idea of “summing all children’s heights” is not expressed with a loop; instead, it is expressed as the additional $HA$ layout field, which has its own computation rule.
@@
A key property of attribute grammar DSLs like this one is that dependencies and traversal orders are static. Specifically, for any field assignment, examining the expression $T$ reveals which other fields on which other nodes it depends on. These static dependencies are critical for invalidation. Also, the schedule language ensures that the relative order in which two fields are computed never changes, even as nodes are added or removed. This is critical for Spineless Traversal. In this paper, we assume that the rules have no cyclic dependencies and that the schedule respects field dependencies; we also assume that the schedule has already been optimized by fusing traversals and ordering field assignments. A substantial literature exists on these topics (grafter; Chen et al., 2022; Liu et al., 2023); the layout implementation in our evaluation is detailed in Section 7.1. In any case, while this paper focuses on web layout, we expect Spineless Traversal to be applicable to incremental computations in other domains, including in compilation, static analysis, computer graphics, and databases, as long as they can be expressed as an attribute grammar in this or a similar DSL.
@@ 3. Incremental Layout
An incremental layout algorithm reuses layout field values between layouts when possible. For example, when the user moves their mouse, a tooltip may need to move to follow the cursor; incremental layout would re-compute the position of the tooltip while reusing all other layout fields for all other elements. Most layouts change only a small fraction of fields—especially the most latency-critical interactions like hovers, drags, animations, and text editing—so incremental layout can be dramatically faster than computing each layout from scratch.